OpenStudy (anonymous):

please help!!!! Input the equation of the given line in standard form. The line including the two points (5, 3), (8, -2)

OpenStudy (anonymous):

SLOPE IS \[\frac{ -2-3 }{ 8-5 }=\frac{ -5 }{ 3 }\]

OpenStudy (anonymous):

\[(y-3)=\frac{ -5 }{ 3 }(x-5)\]

OpenStudy (anonymous):

m = -5/3 and equation will be 3 = -5/3(5)+b => 3 = -25/3+b or b = 34/3 So equation is y = -5/3x+34/3 or 3y = -5x+34 or 5x+3y = 34

OpenStudy (anonymous):

correct at least that is wat i got is what both them got

OpenStudy (anonymous):

what about, The line with m = 3 and b = -2?

OpenStudy (anonymous):

y = 3x-2 or 3x-y = 2

OpenStudy (anonymous):

is that the aswer?

OpenStudy (anonymous):

anyone help with Input the equation of the given line in standard form. The line with m = -1/2 and the point (-2, -3).

OpenStudy (anonymous):

Do you know the equation of a line in standard form?

OpenStudy (anonymous):

It is of the form ax+by+c=0, you can just find the equation of the line as cody did with your previous sum and then convert it to this form. Can you try it?

OpenStudy (anonymous):

no.i just started this but i need help.so cinfused

OpenStudy (anonymous):

Here, you don't have to calculate the slope as it is already given.

OpenStudy (anonymous):

i am still lost:(

OpenStudy (anonymous):